OpenStudy (anonymous):

if y-sin^-1(5x), then dy/dx = ??

OpenStudy (anonymous):

$$y = \sin^{-1}(5x)$$ What is the derivative of \(5x\), and what is the derivative of \(\sin x\)?

OpenStudy (anonymous):

5 and cosx?

OpenStudy (anonymous):

but it is the inverse of sin

OpenStudy (anonymous):

\[\arcsin(x) \neq (\sin(x))^{-1}\]

OpenStudy (anonymous):

\[\frac{ d }{ dx }\arcsin(x) = \frac{ 1 }{ \sqrt{1 - x^2} }\] \[\frac{ d }{ dx }\arcsin(5x) = \frac{ 1 }{ \sqrt{1 - (5x)^2} }\left( \frac{ d }{ dx }(5x) \right) = \frac{ 5 }{ \sqrt{1 - 25x^2} }\]

OpenStudy (anonymous):

ooh, thanks :)

OpenStudy (anonymous):

remember that chain rule is ALWAYS used. except sometimes, you just end up multiplying by one.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i'll remember